# Machine learning for structure-guided materials and process design

### v2

@@ 3.3 Process design using MEG-SGGPO
Experiments are conducted with the following hyperparameters: Deep Q-learning as described in Section 2.1.3 as basic algorithm where the target-network is updated every $n_{\theta}=50$ time-steps. Q-networks with hidden layer sizes of [128, 256, 256, 128], layer normalization and ReLU activation functions. The learning process starts after 100 control-steps. The networks are trained after each control-step with a mini-batch of size 32. The Adam optimizer [cite] is used for neural network training, with a learning rate of $5e^{-4}$. An $\epsilon$-greedy policy, with an initial exploration rate $\epsilon_{0}=0.5$ and the final exploration-rate $\epsilon_{f}=0.0$, with $n_{\epsilon}=390$.
Once a diverse set of crystallographic textures has been identified, MEG-SGGPO is used to guide the metal forming process to the best reachable texture. For solving the identification problem, the MEG-SGGPO approach is allowed to conduct $400$ process runs, so-called episodes. The evolution of the distance between the produced and the chosen goal texture is depicted over the episodes in Fig. 8. Initially, the reinforcement learning agent attempts to identify processing paths that target randomly selected goal textures. As the learning process progresses, the agent focuses increasingly on producing goal texture $\check{g}^{2}$.
Figure 8: Sinkhorn texture distance $\mathscr{D}_{\mathrm{sh}}$ between the produced and the goal texture over the episodes of the MEG-SGGPO approach. The color indicates the targeted texture.
For comparison, Figure 9 presents pole figure plots of the goal texture goal texture $\check{g}^{2}$ and the produced texture $\check{g}^{2^{\prime}}$ after 400 episodes. Visually, the produced texture is highly similar to the goal texture with intensity peaks at the same positions and of a similar magnitude. This also holds for areas not covered by orientations at all. To quantify the effectiveness of the approach, we compute the properties resulting from the produced crystallographic texture and compare them with the desired ones. The results, listed in Table 1, show that five of the properties ($E_{11}$, $E_{22}$, $E_{33}$, $\widetilde{R}_{12}$ and $\widetilde{R}_{13}$) are very close to the target region, while $\widetilde{R}_{23}$ lies inside.
Figure 9: Pole figure plots (001), (110), (111) of the targeted goal texture $\check{g}^{2}$ (top) and the produced texture $\check{g}^{2^{\prime}}$ (bottom)
@@ 4 Discussion
The results presented demonstrate the effectiveness of the SMTLO approach in identifying sets of near-optimal crystallographic textures that exhibit desired properties in a given target region. The set of textures identified by the approach is shown to be more diverse than the benchmark set obtained from the training data set. However, the properties of the identified textures are not completely inside the target region, as can be seen in Fig. 6.
@@
There are two reasons for this: First, the formulation of the optimization objective in the SMTLO approach allows for a trade-off between (i) finding crystallographic textures with properties inside the target region, (ii) identifying possibly diverse textures and, (iii) guaranteeing that texture can be produced by the process under consideration. Second, due to prediction errors of the underlying machine learning model, the SMTLO approach identifies textures whose predicted properties lie inside the target region, while some of the true properties (calculated by the numerical simulation on the basis of the reconstructed textures) slightly lie outside. Both of these issues can be addressed by modifying the objective function of the SMTLO optimizer, defined by Equation 12. For instance, targeting the center of the properties region, instead of its bounds can mitigate both issues. Additionally, the second issue can be addressed by enhancing the machine learning model, for example, by increasing the amount of training data.
@@
Taking a diverse subset of the identified textures as goal textures, the MEG-SGGPO approach guides the metal forming process closely to one of the chosen goal textures. Although the reinforcement learning agent was not able to reproduce one of the goal textures exactly (which is challenging due to the many possible processing paths), the measured distance between the targeted texture $\check{g}^{2}$ and the produced one $\check{g}^{2^{\prime}}$ is sufficiently low ($\mathscr{D}_{\mathrm{sh}}=0.031$) when relating it to the distance to the nearest neighbor in the training data set ($\mathscr{D}_{\mathrm{sh}}=0.033$) and to the furthest data point ($\mathscr{D}_{\mathrm{sh}}=0.141$). We want to remark here, that it took the reinforcement learning agent only 400 episodes to achieve this result compared to the baseline set that is grounded on the generation of 76980 random samples. The reinforcement learning algorithm can therefore be seen as being data efficient.
@@
Nevertheless, in this study, it seems that a lower bound is existing that is difficult to overcome by the MEG-SGGPO approach. In general, this can have two reasons: First, the forming process is unable to produce textures that are closer to the goal texture identified by the SMTLO approach. Yet, the SMTLO approach already addresses this issue by enforcing microstructures to remain within the region delineated by the known microstructures from the training data set via the term $\mathscr{V}_{\mathrm{valid}}\mathscr{C}_{\mathrm{valid}}$ in Equation (12). We expected that a stronger enforcement leads to textures that are better reachable by the process. This, however, can lead to a lower accuracy in terms of obtaining the desired material properties, due to the change in the other weights in the objective function of the SMTLO optimizer. Second, the MEG-SGGPO approach identified a local optimum and got stuck. This can be mitigated by longer MEG-SGGPO runs with optimized hyperparameters. Longer runs, however, have not shown to yield significant improvements in the presented study. As future work, it is desirable to incorporate the knowledge contained in the generated training samples for SMTLO into MEG-SGGPO a priori to enhance its performance.
@@ 5 Conclusion
  • $76980$ samples were sufficient for the SMTLO approach to successfully solve the materials design problem. It allowed us to identify a diverse set of near-optimal textures with desired properties.
  • a total of $400$ episodes was sufficient for the MEG-SGGPO approach to successfully guide the forming process successfully and solve the process design problem.
@@ Acknowledgments
Competing interests
The authors have no competing interests to declare that are relevant to the content of this article.

### v3

@@ 3.3 Process design using MEG-SGGPO
Experiments are conducted with the following hyperparameters: Deep Q-learning as described in Section 2.1.3 as basic algorithm where the target-network is updated every $n_{\theta}=50$ time-steps. Q-networks with hidden layer sizes of [128, 256, 256, 128], layer normalization and ReLU activation functions. The learning process starts after 100 control-steps. The networks are trained after each control-step with a mini-batch of size 32. The Adam optimizer [63] is used for neural network training, with a learning rate of $5\times 10^{-4}$. An $\epsilon$-greedy policy, with an initial exploration rate $\epsilon_{0}=0.5$ and the final exploration-rate $\epsilon_{f}=0.0$, with $n_{\epsilon}=390$.
Once a diverse set of crystallographic textures has been identified, MEG-SGGPO is used to guide the metal forming process to the best reachable crystallographic texture. For solving the identification problem, the MEG-SGGPO approach is allowed to conduct $400$ process runs, so-called episodes. The evolution of the distance between the produced and the chosen goal texture is depicted over the episodes in Fig. 9. Initially, the reinforcement learning agent attempts to identify processing paths that target randomly selected goal textures. As the learning process progresses, the agent focuses increasingly on producing goal texture $\check{g}^{2}$.
Figure 9: Sinkhorn texture distance $\mathscr{D}_{\mathrm{sh}}$ between the produced and the goal texture over the episodes of the MEG-SGGPO approach. The color indicates the targeted crystallographic texture.
For comparison, Figure 10 presents pole figure plots of the goal texture $\check{g}^{2}$ and the produced crystallographic texture $\check{g}^{2^{\prime}}$ after 400 episodes. Visually, the produced crystallographic texture is highly similar to the goal texture with intensity peaks at the same positions and of a similar magnitude. This also holds for areas not covered by orientations at all. To quantify the effectiveness of the approach, we compute the properties resulting from the produced crystallographic texture and compare them with the desired ones. The results, listed in Table 1, show that five of the properties ($E_{11}$, $E_{22}$, $E_{33}$, $\widetilde{R}_{12}$ and $\widetilde{R}_{13}$) are very close to the target region, while $\widetilde{R}_{23}$ lies inside.
Figure 10: Pole figure plots (001), (110), (111) of the targeted goal texture $\check{g}^{2}$ (top) and the produced crystallographic texture $\check{g}^{2^{\prime}}$ (bottom)
@@ 4 Discussion
The results presented demonstrate the effectiveness of the SMTLO approach in identifying sets of near-optimal crystallographic textures that exhibit desired properties in a given target region. The set of crystallographic textures identified by the approach is shown to be more diverse than the benchmark set obtained from the training data set. However, the properties of the identified crystallographic textures are not completely inside the target region, as can be seen in Fig. 7.
@@
There are two reasons for this: First, the formulation of the optimization objective in the SMTLO approach allows for a trade-off between (i) finding crystallographic textures with properties inside the target region, (ii) identifying possibly diverse crystallographic textures and, (iii) guaranteeing that crystallographic texture can be produced by the process under consideration. Second, due to prediction errors of the underlying machine learning model, the SMTLO approach identifies crystallographic textures whose predicted properties lie inside the target region, while some of the true properties (calculated by the numerical simulation on the basis of the reconstructed crystallographic textures) slightly lie outside. Both of these issues can be addressed by modifying the objective function of the SMTLO optimizer, defined by Equation 12. For instance, targeting the center of the properties region, instead of its bounds can mitigate both issues. Additionally, the second issue can be addressed by enhancing the machine learning model, for example, by increasing the amount of training data.
@@
Taking a diverse subset of the identified crystallographic textures as goal textures, the MEG-SGGPO approach guides the metal forming process closely to one of the chosen goal textures. Although the reinforcement learning agent was not able to reproduce one of the goal textures exactly (which is challenging due to the many possible processing paths), the measured distance between the targeted crystallographic texture $\check{g}^{2}$ and the produced one $\check{g}^{2^{\prime}}$ is sufficiently low ($\mathscr{D}_{\mathrm{sh}}=0.031$) when relating it to the distance to the nearest neighbor in the training data set ($\mathscr{D}_{\mathrm{sh}}=0.033$) and to the furthest data point ($\mathscr{D}_{\mathrm{sh}}=0.141$). This can be seen in Figure 9, which depicts the reduction in Sinkhorn distance from the initial texture $g_{0}$ to the final produced texture $\check{g}^{2^{\prime}}$ guided by the reinforcement learning agent.
We want to remark here, that it is generally difficult to evaluate distances between crystallographic textures, even for experts. Therefore, the quality of the produced crystallographic texture is determined by two considerations: (i) the distance of the textures in the microstructure space (as illustrated above), which is statistically valid, and (ii) the distance in properties, which can be assessed based on the results shown in Table 1. It can be seen that for the produced texture, one property lies within the target region, while the remaining five properties are in close proximity. It is remarkable that it took the reinforcement learning agent only 400 episodes to achieve this result compared to the baseline set that is grounded on the generation of 76980 random samples. The reinforcement learning algorithm can therefore be seen as being data efficient.
@@
Nevertheless, in this study, it seems that a lower bound is existing that is difficult to overcome by the MEG-SGGPO approach. In general, this can have two reasons: First, the forming process is unable to produce crystallographic textures that are closer to the goal texture identified by the SMTLO approach. Yet, the SMTLO approach already addresses this issue by enforcing microstructures to remain within the region delineated by the known microstructures from the training data set via the term $\mathscr{V}_{\mathrm{valid}}\mathscr{C}_{\mathrm{valid}}$ in Equation (12). We expected that a stronger enforcement leads to crystallographic textures that are better reachable by the process. This, however, can lead to a lower accuracy in terms of obtaining the desired material properties, due to the change in the other weights in the objective function of the SMTLO optimizer. Second, the MEG-SGGPO approach identified a local optimum and got stuck. This can be mitigated by longer MEG-SGGPO runs with optimized hyperparameters. Longer runs, however, have not shown to yield significant improvements in the presented study. As future work, it is desirable to incorporate the knowledge contained in the generated training samples for SMTLO into MEG-SGGPO a priori to enhance its performance.
@@ 5 Conclusion
  • $76980$ samples were sufficient for the SMTLO approach to successfully solve the materials design problem. It allowed us to identify a diverse set of near-optimal crystallographic textures with desired properties. The crystallographic textures are considered near-optimal in the sense that all of them bear different but satisfactory macroscopic properties and, at the time point of the identification, it is not know which one is best reachable by the process.
  • after solving the materials design problem, a total of $400$ episodes was sufficient for the MEG-SGGPO approach to successfully guide the forming process to the best reachable crystallographic texture and solve the process design problem.
@@ Acknowledgments
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT in order to improve language and readability. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.